find the equation of the line. write the equation in the form Ax+By=C
slope -4/5; passes through (2,4)

Answer by nerdybill(7384)

find the equation of the line. write the equation in the form Ax+By=C
slope -4/5; passes through (2,4)
.
Plug given information into the "point-slope" form:
y - y1 = m(x - x1)
y - 4 = (-4/5)(x - 2)
now, we manipulate the terms until we get the desired form:
begin by multiplying both sides by 5:
5y - 20 = -4(x - 2)
5y - 20 = -4x + 8
5y = -4x + 28
4x + 5y = 28